# UStyle: Waterbody Style Transfer of Underwater Scenes by Depth-Guided Feature Synthesis

### v2

@@ V-E Analysis of Results
Complementary to these numerical results, the Grad-CAM [121] visualizations in Figure 10 illustrate that depth guidance enables more uniform and focused attention across critical image regions, ensuring spatial coherence. These findings suggest that the DA-WCT module guides the network to focus on relevant salient structural cues while mitigating artifacts. Moreover, the training process revealed that UStyle consistently converges across diverse underwater conditions without manual parameter tuning, unlike physics-based models, which are sensitive to reference image quality and scene priors. Overall, UStyle’s design choices yield a robust and efficient framework for underwater STx, and these insights may guide future research in developing even more adaptive and efficient solutions.
@@
Furthermore, during inference using UStyle and the baseline models, we noticed that models without depth guidance, mainly when applied to deep-water style transitions, frequently produced over-saturated outputs with a global averaging effect [122]; this resulted in increased RMSE values by approximately $10\%$ to $12\%$ and LPIPS scores that were consistently higher. In contrast, UStyle maintained stable convergence and demonstrated a more consistent loss decay pattern with a reduction in loss variance of about $15\%$. Furthermore, the inclusion of the CLIP loss [25] in UStyle contributed to a smoother convergence trajectory by aligning semantic features; removal of this component led to slight degradations in SSIM (a reduction of roughly $0.002$ on average) and a marginal increase in RMSE, highlighting the importance of semantic consistency in waterbody STx learning.
@@
Additional observations from our experiments reveal that UStyle’s adaptive multi-stage training strategy [20] resulted in convergence speeds approximately $20\%$ faster than several baseline models. The fine-tuning process on underwater-specific data proved crucial for the network to adapt to domain-specific distortions, yielding significant improvements in performance metrics across diverse underwater scenarios. This indicates that carefully designed loss functions, particularly those focusing on structural similarity [22], color consistency [21], and frequency-domain fidelity [23], are essential for achieving high-quality results in complex underwater environments.
@@
During waterbody fusion, we observed that the depth maps’ quality and depth thresholding precision influenced the stylization process. Variations in depth maps lead to noticeable differences in the retention of fine details, especially in regions with subtle gradient changes. Our future research will emphasize robust depth extraction methods and consider dynamic adjustment of depth-based parameters to enhance performance further. Additionally, the success of UStyle in handling diverse underwater conditions without reliance on manually curated reference images with explicit scene priors suggests that its framework may be extended to other domains where scene priors are challenging to obtain.
Figure 11: Evaluation of UStyle’s generalization capability under three unseen scenarios: (a) Style transfer on previously unseen color hues (yellow and grey) from UIEB [123], (b) Consistent interpolation across progressively intensified waterbody styles from the TURBID [124] dataset, (c) Illumination of low-light underwater scenes from CaveSeg [125] and OceanDark [126] datasets using a well-lit image as style.
@@ VI-A Unseen Style Transfer
The generalization ability of UStyle is validated through three distinct scenarios as shown in Figure 11, using the model trained on the UF7D dataset without any additional fine-tuning. First, in Figure 11 (a) , We used UIEB [123], which feature unseen waterbody hues such as grey and yellow, distinct from the predominantly blue and green hues seen during training. UStyle demonstrated its ability to transfer these unseen hues onto the content images, highlighting its strong generalization capability and robust style representation learned from the UF7D dataset. Next, we examine gradual style intensification using the TURBID [124] dataset in Figure 11 (b). This dataset provides images that are originally clear, along with corresponding versions exhibiting progressively increasing levels of turbidity (light to dark blue tone). Using these as style references, we observe that UStyle performs smooth and consistent style transfer across all levels of distortion, indicating its ability to interpolate within a continuous style manifold.
Figure 11 emphasized the practical significance of UStyle by evaluating style transfer in low-light underwater conditions using images from the CaveSeg [125] and OceanDark [126] datasets. We applied a well-lit, high-quality style image to dimly illuminated underwater scenes. UStyle effectively transferred the brighter illumination conditions from the reference style image, enhancing visibility and clarifying structural details. This capability demonstrates a meaningful practical application, showcasing how UStyle can be effectively utilized for other underwater imaging tasks.
VI-B Computational Analyses
As shown in Table IV, UStyle maintains a stable inference rate across multiple resolutions. Evaluated on a single NVIDIA A100 GPU, UStyle processes a $256\times 256$ image in approximately $518$ ms and a full HD $1920\times 1080$ image in about $741$ ms. While transformer-based methods such as StyTr2 [7] and IEContraAST [5] struggle with high-resolution images due to high memory consumption, UStyle consistently operates across all resolutions, demonstrating scalability. Although MCCNet [115] and ASTMAN [17] achieve faster processing times at low resolutions, their performance is not comparable to UStyle. Moreover, due to memory constraints, StyTr2 [7], ASTMAN [17], and IEContraAST [5] fail to process images at resolutions beyond $1280\times 720$ as well. While UStyle is not intended for real-time applications, rather a high-performance data augmentation and stylazation tool, it still provides competitive processing speeds, especially at high input resolutions.
@@ VI-B Computational Analyses
TABLE IV: Inference speed based on image processing time per image (msec $\downarrow$) across different image resolutions for six data-driven SOTA methods are shown. Evaluations are performed on a single NVidia A100 GPU with $16$ GB RAM; ‘$\times$’ indicates the out-of-memory (OOM) cases.
@@
ASTMAN [17] $180$ $190.8$ $222.7$ $250.6$ $227.8$ $274.0$$\times$
IEContraAST [5] $1219.5$ $2127.7$ $2272.7$ $2702.7$ $3125.0$ $4761.9$$\times$
MCCNet [6] $170.9$ $209.2$ $229.4$ $354.6$ $371.7$ $549.5$ $1886.8$
StyTr2 [7] $136.2$ $274.7$ $285.7$ $1639.3$ $2040.8$$\times$ $\times$
PhotoWCT2 [4] $323.6$ $340.1$ $344.8$ $390.6$ $374.5$ $467.3$ $885.0$
UStyle $518.1$ $621.1$ $641.0$ $645.2$ $645.2$ $740.7$ $1087.0$
VI-C Benefits Over AquaFuse
As we discussed earlier in Sec. IV-C, we use a physics-guided waterbody synthesis module of AquaFuse [1] to extract the waterbody characteristics from the style image. Note that AquaFuse synthesizes the content image by incorporating a reference scene’s waterbody through a physics-guided approximation approach. However, this method relies on a noise-free reference image and prior scene information, including depth, incident angle, and attenuation parameters. As illustrated in Figure 12, AquaFuse is highly sensitive to input parameters such as incident angles and tuned attenuation coefficients, limiting its generalizability across diverse underwater conditions. Its performance is strongly influenced by the choice of reference image and variations in incident angles. In contrast, UStyle, a learning-based STx framework, adapts robustly to varying aquatic environments without requiring manually tuned priors. To facilitate this adaptability, large-scale datasets such as USOD10K and UF7D provide a diverse collection of waterbody styles, enabling more generalized STx learning, as demonstrated in our experimental analyses.
VI-D Limitations and Future Directions
@@ 2. $\bullet$
    Training and inference efficiency: Although the progressive blockwise training strategy enhances convergence stability and multi-scale feature learning, it increases training time compared to single-stage models. Future work could explore the design of lightweight architectures or leverage knowledge distillation techniques [127] to improve efficiency without sacrificing the generalization performance.
@@ 3. $\bullet$
    Extremely turbid or rare waterbody styles: The compounded effects of scattering and absorption can significantly impair style transfer quality for severely turbid scenes [128]. Future work will explore integrating polarization-based enhancement techniques [129] and multi-spectral image enhancement methods [130] to improve UStyle’s robustness in challenging underwater environments. We will also explore adaptive blending mechanisms [131] to better accommodate rare/unseen waterbody styles.
Despite these challenges, UStyle significantly advances underwater STx literature for data-driven waterbody fusion. Future research will optimize its efficiency, extend it to real-time video applications, and improve robustness in complex underwater scenarios.
@@ VII Conclusion
In this work, we presented UStyle, the first data-driven framework for underwater waterbody style transfer that robustly fuses waterbody characteristics without relying on manually curated reference images with explicit scene priors. Leveraging a ResNet-based encoder-decoder architecture, a progressive blockwise training strategy, and a novel depth-aware whitening and coloring transform (DA-WCT), UStyle effectively preserves structural integrity while achieving perceptually compelling stylization. Extensive quantitative and qualitative evaluations demonstrate that UStyle outperforms SOTA artistic, photorealistic, and physics-based methods, achieving improvements of up to $16\%$ in RMSE, reaching PSNR values of $29.14$  dB, SSIM values as high as $0.9613$, and attaining LPIPS scores as low as $0.1373$. Ablation studies further confirm that integrating depth supervision and additional regularization losses is critical for enhancing geometric consistency and minimizing perceptual distortions. Moreover, computational analyses reveal that UStyle maintains stable inference rates across various resolutions, showcasing its scalability for high-resolution underwater imaging applications. Although challenges remain in handling extreme turbidity and further optimizing training efficiency, our findings underscore the potential of UStyle as a robust tool for underwater imaging, data augmentation, and vision tasks. Future work will incorporate depth estimation, explore lightweight architectures for real-time applications, and extend the framework to handle more complex underwater STx scenarios in real-time.
Acknowledgments

### v3

@@ V-E Analysis of Results
Complementary to these numerical results, the Grad-CAM [selvaraju2017grad] visualizations in Fig. 8 illustrate that depth guidance enables more uniform and focused attention across critical image regions, ensuring spatial coherence. These findings suggest that the DA-WCT module guides the network to focus on relevant salient structural cues while mitigating artifacts. Moreover, the training process revealed that UStyle consistently converges across diverse underwater conditions without manual parameter tuning, unlike physics-based models, which are sensitive to reference image quality and scene priors.
@@
During inference using UStyle and the baseline models, we noticed that models without depth guidance, mainly when applied to deep-water style transitions, frequently produced over-saturated outputs with a global averaging effect [reinhard2001color]; this resulted in increased RMSE values by approximately $10\%$ to $12\%$ and LPIPS scores that were consistently higher. In contrast, UStyle maintained stable convergence and demonstrated a more consistent loss decay pattern with a reduction in loss variance of about $15\%$. Furthermore, the inclusion of the CLIP loss [radford2021learning] in UStyle contributed to a smoother convergence trajectory by aligning semantic features; removal of this component led to slight degradations in SSIM (a reduction of roughly $0.002$ on average) and a marginal increase in RMSE, highlighting the importance of semantic consistency in waterbody STx learning.
@@
Additional observations from our experiments reveal that UStyle’s adaptive multi-stage training strategy [wang2018progressive] resulted in convergence speeds approximately $20\%$ faster than several baseline models. The fine-tuning process on underwater-specific data proved crucial for the network to adapt to domain-specific distortions, yielding significant improvements in performance metrics across diverse underwater scenarios. This indicates that carefully designed loss functions, particularly those focusing on structural similarity [hore2010image], color consistency [dong2022underwater], and frequency-domain fidelity [brigham1988fast], are essential for achieving high-quality results in complex underwater environments.
@@
Furthermore, as shown in Table IV, noise in the estimated depth maps and the precision of depth thresholding significantly influence the STx performance in UStyle. Noisy depth maps lead to spatial imbalance in the style transfer, causing uneven blending and reduced retention of fine texture details. In particular, regions with subtle depth gradients are highly sensitive to noise, where Monodepth2 introduces artifacts and inconsistent stylization, while more accurate estimators such as DepthAnythingV2 preserve smooth transitions and maintain structural fidelity in the stylized outputs. Our future research will emphasize robust depth extraction methods and consider dynamic adjustment of depth-based parameters to enhance performance further. Additionally, the success of UStyle in handling diverse underwater conditions without reliance on manually curated reference images with explicit scene priors suggests that its framework may be extended to other domains where scene priors are challenging to obtain.
@@ VI-A Unseen Style Transfer
The generalization capability of UStyle is evaluated under three distinct scenarios as illustrated in Fig. 9, using the model trained on the UF7D dataset without any additional fine-tuning. In the first scenario, UIEB dataset [UIEB_li] is used to assess style transfer to unseen waterbody hues, such as grey and yellow, which differ from the predominantly blue and green tones present in the training data. As Fig. 9(a) shows, UStyle successfully transfers these previously unseen hues onto the content images, demonstrating its strong generalization and robust style representation learned from the UF7D.
In the second scenario, TURBID dataset [duarte2016dataset] is used to examine gradual style intensification. This dataset provides clear underwater images alongside corresponding versions exhibiting progressively increasing turbidity, ranging from light to dark blue tones. When these were used as style references, UStyle achieves smooth and consistent STx across all levels of turbidity, indicating its ability to interpolate within a continuous style manifold. Finally, Fig. 9(c) highlights the practical applicability under low-light underwater conditions, evaluated using CaveSeg [abdullah2023caveseg] and OceanDark [Oceandark_Marques_2020_CVPR_Workshops] datasets. For validation, we apply a well-lit, high-quality reference style image to dimly illuminated underwater scenes. As the results show, UStyle effectively transfers the brighter illumination characteristics, enhancing visibility and preserving structural details. These results demonstrate UStyle’s potential for practical underwater image enhancement, beyond stylization alone.
Figure 9: Evaluation of UStyle’s generalization capability under three unseen scenarios: (a) Style transfer on previously unseen color hues (yellow and grey) from UIEB [UIEB_li]; (b) Consistent interpolation across progressively intensified waterbody styles from TURBID [duarte2016dataset] dataset; and (c) Illumination of low-light underwater scenes from CaveSeg [abdullah2023caveseg] and OceanDark [Oceandark_Marques_2020_CVPR_Workshops] datasets using a well-lit image as style.
VI-B Benefits Over AquaFuse
As we discussed earlier in Sec. IV-C, we use a physics-guided waterbody synthesis module of AquaFuse [siddique2024aquafuse] to extract the waterbody characteristics from the style image. Note that AquaFuse synthesizes the content image by incorporating a reference scene’s waterbody through a physics-guided approximation approach. However, this method relies on a noise-free reference image and prior scene information, including depth, incident angle, and attenuation parameters. As illustrated in Fig. 10, AquaFuse is highly sensitive to input parameters such as incident angles and tuned attenuation coefficients, limiting its generalizability across diverse underwater conditions. Its performance is strongly influenced by the choice of reference image and variations in incident angles. In contrast, UStyle, a learning-based STx framework, adapts robustly to varying aquatic environments without requiring manually tuned priors. To facilitate this adaptability, large-scale datasets such as USOD10K and UF7D provide a diverse collection of waterbody styles, enabling more generalized STx learning, as demonstrated in our experimental analyses.
@@ VI-B Benefits Over AquaFuse
TABLE V: Inference speed based on image processing time per image (msec $\downarrow$) across different image resolutions for six data-driven SOTA methods are shown. Evaluations are performed on a single NVidia A100 GPU with $16$ GB RAM; $\times$’ indicates the out-of-memory (OOM) cases.
@@
ASTMAN [deng2020arbitrary] $180$ $190.8$ $222.7$ $250.6$ $227.8$ $274.0$ $\times$
IEContraAST [chen2021artistic] $1219.5$ $2127.7$ $2272.7$ $2702.7$ $3125.0$ $4761.9$ $\times$
MCCNet [deng2021arbitrary] $170.9$ $209.2$ $229.4$ $354.6$ $371.7$ $549.5$ $1886.8$
StyTr2 [deng2022stytr2] $136.2$ $274.7$ $285.7$ $1639.3$ $2040.8$ $\times$ $\times$
PhotoWCT2 [chiu2022photowct2] $323.6$ $340.1$ $344.8$ $390.6$ $374.5$ $467.3$ $885.0$
UStyle (ours) $518.1$ $621.1$ $641.0$ $645.2$ $645.2$ $740.7$ $1087.0$
VI-C Computational Analyses
As shown in Table V, UStyle maintains a stable inference rate across multiple resolutions. Evaluated on a single NVIDIA A100 GPU, UStyle processes a $256\times 256$ image in approximately $518$ ms and a full HD $1920\times 1080$ image in about $741$ ms. While transformer-based methods such as StyTr2 [deng2022stytr2] and IEContraAST [chen2021artistic] struggle with high-resolution images due to high memory consumption, UStyle consistently operates across all resolutions, demonstrating scalability. Although MCCNet [deng2021stytr2] and ASTMAN [deng2020arbitrary] achieve faster processing times at low resolutions, their performance is not comparable to UStyle. Moreover, due to memory constraints, StyTr2 [deng2022stytr2], ASTMAN [deng2020arbitrary], and IEContraAST [chen2021artistic] fail to process images at resolutions beyond $1280\times 720$ as well. While UStyle is not intended for real-time applications, rather a high-performance data augmentation and stylazation tool, it still provides competitive processing speeds, especially at high input resolutions.
Figure 11: Qualitative comparison of underwater STx methods on the UF7D-Challenge dataset. The top row shows different waterbody styles, and each column applies one style to distinct content images. Subsequent rows present outputs from three top-performing models: IEContrAST, PhotoWCT2, and UStyle. As shown, UStyle outperforms the other models and handles the illumination, color tones, and scene structure to generate more realistic underwater stylization.
TABLE VI: Quantitative comparison of PSNR ($\uparrow$), SSIM ($\uparrow$) and RMSE ($\downarrow$) for UStyle and two best-performing STx methods: IEContrAST and PhotoWCT2. Averaged scores on $100$ ($10\times 10$ combinations) for each case in UF7D-Challenge data are reported; best scores are highlighted in bold.
Metric Model B $\rightarrow$ G G $\rightarrow$ B DG $\rightarrow$ DB DB $\rightarrow$ DG DB $\rightarrow$ C DG $\rightarrow$ C BG $\rightarrow$ C GB $\rightarrow$ C
PSNR IEContrAST 15.58 14.83 17.32 16.99 12.57 14.013 14.41 14.47
PhotoWCT2 16.39 16.37 18.05 18.80 13.69 13.59 15.69 14.69
UStyle (ours) 21.21 22.15 25.55 23.31 20.30 20.99 20.47 20.98
PSNR IEContrAST 0.4522 0.4829 0.5811 0.5510 0.4066 0.4680 0.4522 0.4718
PhotoWCT2 0.7479 0.7814 0.7090 0.7454 0.6534 0.6418 0.7641 0.6990
UStyle (ours) 0.8758 0.9047 0.8448 0.8544 0.8695 0.8922 0.8733 0.8822
RMSE IEContrAST 43.83 49.24 36.45 37.87 63.67 53.71 50.16 50.40
PhotoWCT2 43.46 42.86 35.59 32.62 60.90 60.16 45.32 52.62
UStyle (ours) 25.45 23.13 18.30 21.05 28.54 26.91 27.19 25.84
VI-D Generalization on Uf7D-Challenge
To further evaluate the generalization performance of UStyle, we curate the UF7D-Challenge test set, which contains $70$ images with $10$ samples for each of the seven waterbody styles in UF7D. The dataset spans a broad range of environments, including open-ocean scenes, underwater caves, and indoor swimming pools, capturing variations in water clarity, color, and structural complexity. It also encompasses diverse illumination conditions: from low-light environments to brightly lit shallow-water scenes – reflecting natural temporal and lighting variations. The images originate from both online open-access sources and in-field captures using various underwater cameras. Although the exact sensor specifications are unknown, the dataset naturally captures cross-sensor variability due to differences in optics, color calibration, and acquisition settings. Overall, UF7D-Challenge provides a benchmark for evaluating underwater style transfer under heterogeneous environmental and sensor conditions.
[…]
VI-E Limitations and Future Directions
@@ 2. $\bullet$
    Training and inference efficiency: Although the progressive blockwise training strategy enhances convergence stability and multi-scale feature learning, it increases training time compared to single-stage models. Future work could explore the design of lightweight architectures or leverage knowledge distillation techniques [hinton2015distilling] to improve efficiency without sacrificing the generalization performance.
@@ 3. $\bullet$
    Extremely turbid or rare waterbody styles: The compounded effects of scattering and absorption can significantly impair style transfer quality for severely turbid scenes [ancuti2012enhancing]. techniques. Future work will explore integrating polarization-based enhancement [schechner2006recovery] and multi-spectral image enhancement methods [aggarwal2020fusion] to improve UStyle’s robustness in challenging underwater environments. We will also explore adaptive blending mechanisms [park2019semantic] to better accommodate rare/unseen waterbody styles.
Despite these challenges, UStyle significantly advances underwater STx literature for data-driven waterbody fusion. Future research will optimize efficiency and extend it to real-time video applications.
@@ VII CONCLUSION
In this work, we presented UStyle, the first data-driven underwater waterbody style transfer framework that fuses waterbody characteristics without curated references or explicit scene priors. Leveraging a ResNet-based encoder-decoder, progressive blockwise training, and a novel depth-aware whitening and coloring transform (DA-WCT), UStyle preserves structural integrity while achieving compelling stylization. Extensive evaluations show that UStyle outperforms SOTA artistic, photorealistic, and physics-based methods – with up to $16\%$ RMSE improvement, PSNR of $29$ dB, SSIM of $0.96$, and LPIPS as low as $0.13$. Ablation studies confirm that depth supervision and regularization losses enhance geometric consistency and reduce distortions. Moreover, computational analyses reveal that UStyle maintains stable inference rates across various resolutions, showcasing its scalability for high-resolution underwater imaging applications. Although challenges remain in handling extreme turbidity and further optimizing training efficiency, UStyle shows promise for underwater imaging, data augmentation, and vision tasks. Future work will focus on improved depth estimation, lightweight real-time architectures, and extending the framework to more complex scenarios.
ACKNOWLEDGMENT